OpenStudy (anonymous):

the perimeter of a rectangle is 18 feet, and the area of the rectangle is 20 square feet. What is the length of the rectangle?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

cuz 4 x 5 = 20 (area) and 5+5=10, 4+4=8, 10+8=18 :) (perimeter